OpenStudy (anonymous):

Write the equation of the quadratic function with roots -1 and 1 and a vertex at (0, -3).

OpenStudy (triciaal):

the roots of a quadratic occur when (x-a) is a factor and f(x) = 0 ; a is the root f(x) = (x+1)(x-1)

OpenStudy (triciaal):

general form for a quadratic ax^2 + bx + c =0 vertex form a(x-h)^2 + k = 0 a=a. h = -b/2a and k = f(h)

OpenStudy (triciaal):

 I'm stuck on a simple problem!